OpenStudy (anonymous):

Factor. 2x2 – x – 6 A. (2x + 3)(x – 2) B. (2x – 3)(x – 2) C. (2x – 2)(x + 3) D. (2x – 3)(x + 2)

OpenStudy (anonymous):

(2x+3)(x−2)

OpenStudy (jadeishere):

(2x - 3)(x + 2)

OpenStudy (jadeishere):

-3x + 2x = -x 3x - 2x = +x

OpenStudy (anonymous):

so A?

OpenStudy (anonymous):

yupperdoodles :)

OpenStudy (anonymous):

Factor completely. 3x3 – 48x A. 3x(x – 16) B. 3x(x – 4)(x – 4) C. 3(x + 4)(x – 4) D. 3x(x – 4)(x + 4)

OpenStudy (anonymous):

@DemolisionWolf

OpenStudy (anonymous):

3x(x+4)(x−4)

OpenStudy (anonymous):

It's C :)

OpenStudy (anonymous):

3x3 – 48x first, factor out all common factors from both terms, we have an 'x' in both and... a 3 48 = 3 * 16 3x( x^2 - 16) recognize that 16 is a square, its 4^2 re write the expression (3x) (x^2 - 4^2) looks like we have a 'difference of squres' they look like this: (A^2 - B^2) = (A - B) (A + B) can you rewrite our expressing using this method of 'differnce of squres'?